OpenStudy (anonymous):

Check for GCF then factor completely by grouping 2x^3 + 4x^2 - 16x -32

OpenStudy (anonymous):

\[2x^3 + 4x^2 - 16x -32=2(x^3+2x^2-8x-16)\]\[=2\left[ x^2(x+2)-8(x+2) \right]\]\[=2(x+2)(x^2-8)\]